# CALT-TH 2023-050

### v1

@@ Abstract
We show via explicit construction that all the extreme rays of both the 3-party quantum entropy cone and the 4-party stabilizer entropy cone can be obtained from subsystem coarse-grainings of specific higher-party quantum states, namely extreme states characterized by saturating a (non-trivial) maximal set of instances of subadditivity. This suggests that the study of the “subadditivity cone”, and the set of its extreme rays realizable in quantum mechanics, provides a powerful new tool for deriving inner bounds for the quantum and stabilizer entropy cones, as well as constraints on new inequalities for the von Neumann entropy.
@@ I Introduction
The derivation of the fundamental inequalities satisfied by the von Neumann entropy is an important problem in quantum information theory. Unfortunately, it is also notoriously difficult. Since the proof of strong subadditivity by Lieb and Ruskai half a century ago Lieb and Ruskai (1973a, b), no new unconstrained inequality has been found. New inequalities involving four or more parties were discovered in Linden and Winter (2005); Cadney et al. (2012), but they are constrained, meaning that they only apply to particular density matrices which saturate other inequalities. In the classical case, i.e., for the Shannon entropy, new inequalities have been derived in Zhang and Yeung (1998); Makarychev et al. (2002); Zhang (2003); Matús (2007), and Cadney et al. (2012) speculated that they might hold also in quantum mechanics. However, to our knowledge, at present there is no strong argument that new unconstrained inequalities for the von Neumann entropy should exist.
The main goal of this letter is to show that a seemingly much simpler problem, the quantum marginal independence problem (QMIP) Hernández-Cuenca et al. (2019), originally motivated by an approach Hubeny et al. (2018, 2019); Hernández-Cuenca et al. (2022) to the study of entropy inequalities in quantum gravity Ryu and Takayanagi (2006); Hayden et al. (2013); Bao et al. (2015), might provide a powerful tool for the derivation of constraints on the existence of new inequalities and their structure. Specifically, we will show that already from a partial solution to the “extremal” version of the QMIP up to nine parties, we can derive definite inner bounds for the quantum entropy cone, which in the case of three parties coincide with the full cone Pippenger (2003), and in the case of four parties coincide with the entropy cone of stabilizer states Linden et al. (2013).
The structure of the paper is as follows. In §II and §III, we review the main definitions concerning entropy cones, the QMIP, and its extremal version. In §IV, we formulate the inner bound to the quantum entropy cone for an arbitrary number of parties. In §V, we review the definition of hypergraph models from Bao et al. (2020a) and the main result from Walter and Witteveen (2021) about their realizability by stabilizer states. We then use this technology in §VI to derive the main result of this letter, and conclude in §VII with a list of open questions.
@@ II The quantum entropy cone
For any ${\sf N}$, an outer bound of $\overline{\Sigma}^{*}_{{\sf N}}$ is given by the polyhedral cone obtained via the intersection of the half-spaces specified by all instances of subadditivity (SA) and strong subadditivity (SSA) of the von Neumann entropy, and we will denote this cone by $\Sigma_{{\sf N}}$. For ${\sf N}\leq 3$ it is easy to show that $\overline{\Sigma}^{*}_{{\sf N}}=\Sigma_{{\sf N}}$ by constructing quantum states that realize entropy vectors belonging to the extreme rays of $\Sigma_{{\sf N}}$. On the other hand, for ${\sf N}\geq 4$, the constrained inequalities found in Linden and Winter (2005); Cadney et al. (2012) imply that this construction is no longer possible, since there are regions on the boundary of $\Sigma_{{\sf N}}$ which are inaccessible to quantum states. This however is not enough to conclude that $\overline{\Sigma}^{*}_{{\sf N}}\subset\Sigma_{{\sf N}}$ (strictly), since there might be quantum states that approximate the entropy vectors in these regions arbitrarily well.
@@ III Quantum marginal independence
An interesting and well-known problem in quantum mechanics, where entropy inequalities play a central role, is the quantum marginal problem Schilling (2014). Given density matrices $\rho_{\mathscr{I}}$ for some subsystems, it asks whether there exists a global density matrix $\rho$ such that all the given $\rho_{\mathscr{I}}$ can be obtained from $\rho$ as marginals. Entropy inequalities restrict the set-ups for which a solution exists.
@@
The ${\sf N}$-party subadditivity cone (SAC) is defined as the polyhedral cone in entropy space carved out by all instances of SA for that ${\sf N}$.
Consider a face $\mathscr{F}$ of the SAC and a vector $\vec{\sf S}\in\text{int}(\mathscr{F})$. Notice that the collection of SA instances which are saturated by $\vec{\sf S}$ is independent from the specific choice of this vector. The saturation of SA is equivalent to the vanishing of the mutual information, which occurs if and only two subsystems are independent. We can then interpret $\text{int}(\mathscr{F})$ to correspond to a specification of which subsystems are correlated and which are independent, while remaining agnostic about the specific values of the entropies. In its original formulation Hernández-Cuenca et al. (2019), the QMIP asked for which faces does $\text{int}(\mathscr{F})$ contain at least one entropy vector that can be realized by a quantum state. We now slightly generalize this question by replacing “realized” with “approximated arbitrarily well.” In this generalized version, the QMIP can then be trivially solved if $\overline{\Sigma}^{*}_{{\sf N}}$ is known.
In this letter we proceed in the opposite direction and attempt to extract information about $\overline{\Sigma}^{*}_{{\sf N}}$ from the solution to the QMIP for different values of ${\sf N}^{\prime}\geq{\sf N}$. In particular, we will be interested in the extremal version of the QMIP, which we denote as EQMIP, where we only focus on the 1-dimensional faces of the SAC, i.e., its extreme rays. Denoting by ^$\widehat{\mathscr{R}}_{{\sf N}}$ the set of all extreme rays of the ${\sf N}$-party SAC, the solution $\mathscr{R}_{{\sf N}}$ to the EQMIP is then
$\mathscr{R}_{{\sf N}}:=\widehat{\mathscr{R}}_{{\sf N}}\cap\overline{\Sigma}^{*}_{{\sf N}}$ (2)
@@ IV Inner bounds from extremal marginal independence
IV.1 Subsystem coarse-grainings
Given a density matrix $\rho_{{\sf N}^{\prime}}$, we want to consider a coarse-graining of the ${\sf N}^{\prime}$ parties into ${\sf N}$ composite ones and to relate the entropy vectors before and after the coarse-graining. We will also consider purifications of $\rho_{{\sf N}^{\prime}}$, and allow for coarse-grainings of the “purifying” party. Defining $\llbracket{\sf N}\rrbracket=\{0,1,\ldots,{\sf N}\}$, where we added 00 to account for the purifier, such a coarse-graining can then be specified by a surjective function
@@ IV.1 Subsystem coarse-grainings
which specifies, for each party $\ell^{\prime}\in\llbracket{\sf N}^{\prime}\rrbracket$, which of the coarse-grained $\llbracket{\sf N}\rrbracket$ parties it belongs to.
Given a density matrix $\rho_{{\sf N}^{\prime}}$ with entropy vector $\vec{{\sf S}}^{\prime}$, and a coarse-graining $f$, the components of the entropy vector of the same density matrix after the coarse graining are then given by
${\sf S}_{\mathscr{I}}={\sf S}_{f^{-1}(\mathscr{I})},$ (4)
@@
Notice that in general, this map of entropy vectors from $\Sigma_{{\sf N}^{\prime}}^{*}$ to $\Sigma_{{\sf N}}^{*}$, which we denote by $\Psi_{f}$, can formally be extended to vectors in the full $\overline{\Sigma}^{*}_{{\sf N}^{\prime}}$, and can map a vector on the boundary of $\overline{\Sigma}^{*}_{{\sf N}^{\prime}}$ to one strictly in the interior of $\overline{\Sigma}^{*}_{{\sf N}}$. For example, any coarse-graining to ${\sf N}=2$ of the ${\sf N}=3$ entropy vector corresponding to the 4-party GHZ state gives an entropy vector strictly inside $\overline{\Sigma}^{*}_{2}$ (since no instance of SA is saturated).
@@ IV.2 Inner bounds for $\overline{\Sigma}^{*}_{{\sf N}}$
Using this construction, we can now derive an inner bound for $\overline{\Sigma}^{*}_{{\sf N}}$ as follows. Suppose that $\mathscr{R}_{{\sf N}^{\prime}}$ is known for some ${\sf N}^{\prime}\geq{\sf N}$. We can then consider all possible coarse-graining from ${\sf N}^{\prime}$ to ${\sf N}$, mapping all vectors in $\mathscr{R}_{{\sf N}^{\prime}}$ to ${\sf N}$-party entropy vectors, and taking the conical hull. Formally, for any given ${\sf N}$ and ${\sf N}^{\prime}\geq{\sf N}$, we define
@@
It suffices to show that for any ${\sf N}^{\prime}\geq{\sf N}$ and $\vec{\sf R}\in\mathscr{R}_{{\sf N}}$, there exists some $\vec{\sf R}^{\prime}\in\mathscr{R}_{{\sf N}^{\prime}}$ and a coarse-graining $f$ from ${\sf N}^{\prime}$ to ${\sf N}$ such that $\Psi_{f}(\vec{\sf R}^{\prime})=\vec{\sf R}$. For the case where $\vec{\sf R}$ is realized by a density matrix $\rho$, we can realize the desired $\vec{\sf R}^{\prime}$ with $\rho^{\prime}=\rho\otimes\ket{{\bf 0}}\!\bra{{\bf 0}}_{[{\sf N}^{\prime}]\setminus[{\sf N}]}$, and choose $f$ such that the additional factors are coarse-grained with the purifier for $\rho^{\prime}$. For the general case, we refer the reader to Section 3.4 of He et al. (2023). ∎
@@ V Hypergraph models
The prescription we just described associates to each ${\sf N}$-party hypergraph model a vector in $\mathbb{R}^{\sf{D}}$. The collection of all these vectors (at fixed ${\sf N}$) is again a convex cone  [^3] A conical combination of two entropy vectors realized by hypergraphs is realized by the hypergraph obtained by the disjoint union of the hypergraphs with the weights appropriately rescaled., and we have Walter and Witteveen (2021):
@@
While in principle hypergraph models are not sufficient to solve the EQMIP completely at arbitrary ${\sf N}$ (see §VII), they at least provide a partial solution that will be sufficient for our purposes.
@@
Figure 1: A hypergraph giving an element of $\mathscr{R}_{6}$. The hyperedge (yellow) has weight 2, and edges (blue) have weight 1.
@@
Figure 2: A hypergraph giving an element of $\mathscr{R}_{9}$. The hyperedge (yellow) has weight 2, and edges (blue) have weight 1.
@@
Figure 3: A hypergraph giving an element of $\mathscr{R}_{8}$. The two hyperedges (yellow and purple) have weight 2, and edges (blue) have weights either 1 (when unspecified) or 2.
@@ VI Constructions for small ${\sf N}$
Having reviewed the necessary tools from hypergraph models, we now use this technology to construct inner bounds for $\overline{\Sigma}^{*}_{{\sf N}}$ for ${\sf N}\leq 4$. The logic will be as follows: we will construct hypergraphs that realize certain extreme rays of the SAC which, by 1, are automatically solutions to the EQMIP. Even without a full solution to the EQMIP, we will then show via particular coarse-grainings that we can obtain entropy vectors for a smaller number of parties, which give powerful inner bounds.
@@ VI.1 Bounds for ${\sf N}=2$
This case it trivial, but we include it for completeness. The outer bound $\Sigma_{2}$ to $\overline{\Sigma}^{*}_{2}$ is the SAC, and its extreme rays are realized by Bell pairs. We then have
@@ VI.2 Bounds for ${\sf N}=3$
As we mentioned in §II, it was already shown in Pippenger (2003) that $\overline{\Sigma}^{*}_{3}=\Sigma_{3}$. This equivalence follows from the fact that the extreme rays of $\Sigma_{3}$ contain entropy vectors realized by Bell pairs, the 4-party GHZ state, and the 4-party absolutely maximally entangled state (also known as the 4-party perfect tensor) Helwig and Cui (2013). With the only exception of the 4-party GHZ, all these states also generate extreme rays of the 3-party SAC and are hence elements of $\mathscr{R}_{3}$. Their conical hull is $\Delta_{3}^{3}$, which coincides with the 3-party holographic entropy cone Bao et al. (2015).
To show that the sequence in Eq. (6) converges to $\overline{\Sigma}^{*}_{3}$, we need to consider $\Delta_{3}^{3+i}$ for $i\geq 1$. The sets $\mathscr{R}_{4}$ and $\mathscr{R}_{5}$ were already found in Hernández-Cuenca et al. (2019), and it is straightforward to verify that they imply $\Delta_{3}^{3}=\Delta_{3}^{4}=\Delta_{3}^{5}$. Hence, we need $i\geq 3$. It turns out that $i=3$ is sufficient, since one can verify that the entropy vector associated to the hypergraph in Figure 1 is in $\mathscr{R}_{6}$, and under the coarse-graining

### v2

@@ Abstract
We show via explicit construction that all the extreme rays of both the three-party quantum entropy cone and the four-party stabilizer entropy cone can be obtained from subsystem coarse grainings of specific higher-party quantum states, namely extreme states characterized by saturating a (non-trivial) maximal set of instances of subadditivity. This suggests that the study of the “subadditivity cone”, and the set of its extreme rays realizable in quantum mechanics, provides a powerful tool for deriving inner bounds for the quantum and stabilizer entropy cones, as well as constraints on new inequalities for the von Neumann entropy.
@@ I Introduction
The derivation of the fundamental inequalities satisfied by the von Neumann entropy is an important problem in quantum information theory. Unfortunately, it is also notoriously difficult. Since the proof of strong subadditivity (SSA) by Lieb and Ruskai half a century ago Lieb and Ruskai (1973a, b), no new unconstrained inequality has been found. New inequalities involving four or more parties were discovered in Linden and Winter (2005); Cadney et al. (2012); Christandl et al. (2023), but they are constrained, meaning that they only apply to particular density matrices which saturate other entropic constraints. In the classical case, i.e., for the Shannon entropy, new inequalities have been derived in Zhang and Yeung (1998); Makarychev et al. (2002); Zhang (2003); Matús (2007), and Cadney et al. (2012) speculated that they might hold also in quantum mechanics. However, to our knowledge, at present there is no strong argument that new unconstrained inequalities for the von Neumann entropy should exist.
The main goal of this paper is to show that a seemingly much simpler problem, the quantum marginal independence problem (QMIP) Hernández-Cuenca et al. (2019), originally motivated by an approach Hubeny et al. (2018, 2019); Hernández-Cuenca et al. (2022) to the study of entropy inequalities in quantum gravity Ryu and Takayanagi (2006); Hayden et al. (2013); Bao et al. (2015), might provide a powerful tool for the derivation of constraints on the existence of new inequalities and their structure. Specifically, we will show that already from a partial solution to the “extremal” version of the QMIP up to nine parties, we can derive definite inner bounds for the quantum entropy cone, which in the case of three parties coincide with the full cone Pippenger (2003), and in the case of four parties coincide with the entropy cone of stabilizer states Linden et al. (2013). We stress that even though we are not in this paper deriving new bounds on quantum entropy cones for four or more parties, our method introduces a conceptual idea that may prove fruitful to explore further.
The structure of the paper is as follows. In §II and §III, we review the main definitions concerning entropy cones, the QMIP, and its extremal version. In §IV, we formulate the inner bound to the quantum entropy cone for an arbitrary number of parties. In §V, we review the definition of hypergraph models from Bao et al. (2020a) and the main result from Walter and Witteveen (2021) about their realizability by stabilizer states. We then use this technology in §VI to derive the main result of this paper, and conclude in §VII with a list of open questions.
@@ II The quantum entropy cone
For any ${\sf N}$, an outer bound of $\overline{\Sigma}^{*}_{{\sf N}}$ is given by the polyhedral cone obtained via the intersection of the half-spaces specified by all instances of subadditivity (SA) and strong subadditivity (SSA) of the von Neumann entropy, and we will denote this cone by $\Sigma_{{\sf N}}$ Pippenger (2003). For ${\sf N}\leq 3$ it is easy to show that $\overline{\Sigma}^{*}_{{\sf N}}=\Sigma_{{\sf N}}$ by constructing quantum states that realize entropy vectors belonging to the extreme rays of $\Sigma_{{\sf N}}$ Pippenger (2003). On the other hand, for ${\sf N}\geq 4$, the constrained inequalities found in Linden and Winter (2005); Cadney et al. (2012) imply that this construction is no longer possible, since there are regions on the boundary of $\Sigma_{{\sf N}}$ which are inaccessible to quantum states. This however is not enough to conclude that $\overline{\Sigma}^{*}_{{\sf N}}\subset\Sigma_{{\sf N}}$ (strictly), since there might be quantum states that approximate the entropy vectors in these regions arbitrarily well.
@@ III Quantum marginal independence
An interesting and well-known problem in quantum mechanics is the quantum marginal problem Schilling (2014). Given density matrices $\rho_{\mathscr{I}}$ for some subsystems, it asks whether there exists a global density matrix $\rho$ such that all the given $\rho_{\mathscr{I}}$ can be obtained from $\rho$ as marginals. In this context, entropy inequalities provide necessary conditions for the existence of a solution Carlen et al. (2013).
@@
The ${\sf N}$-party subadditivity cone (SAC) is defined as the polyhedral cone in entropy space carved out by all instances of SA for that ${\sf N}$. Consider a face $\mathscr{F}$ of the SAC and a vector $\vec{\sf S}\in\text{int}(\mathscr{F})$. Notice that the collection of SA instances which are saturated by $\vec{\sf S}$ is independent from the specific choice of this vector. The saturation of SA is equivalent to the vanishing of the mutual information, which occurs if and only two subsystems are independent. We can then interpret $\text{int}(\mathscr{F})$ to correspond to a specification of which subsystems are correlated and which are independent, while remaining agnostic about the specific values of the entropies. In its original formulation Hernández-Cuenca et al. (2019), the QMIP asked for which faces does $\text{int}(\mathscr{F})$ contain at least one entropy vector that can be realized by a quantum state. We now slightly generalize this question by replacing “realized” with “approximated arbitrarily well.” In this generalized version, the QMIP can then be trivially solved if $\overline{\Sigma}^{*}_{{\sf N}}$ is known.
In this paper we proceed in the opposite direction and attempt to extract information about $\overline{\Sigma}^{*}_{{\sf N}}$ from the solution to the QMIP for different values of ${\sf N}^{\prime}\geq{\sf N}$. In particular, we will be interested in the extremal version of the QMIP, which we denote as EQMIP, where we only focus on the one-dimensional faces of the SAC, i.e., its extreme rays. Denoting by ^$\widehat{\mathscr{R}}_{{\sf N}}$ the set of all extreme rays of the ${\sf N}$-party SAC, the solution $\mathscr{R}_{{\sf N}}$ to the EQMIP is then
$\mathscr{R}_{{\sf N}}:=\widehat{\mathscr{R}}_{{\sf N}}\cap\overline{\Sigma}^{*}_{{\sf N}}.$ (2)
@@ IV Inner bounds from extremal marginal independence
IV.1 Subsystem coarse grainings
Given a density matrix $\rho_{{\sf N}^{\prime}}$, we want to consider a coarse graining of the ${\sf N}^{\prime}$ parties into ${\sf N}$ composite ones and to relate the entropy vectors before and after the coarse graining. We will also consider purifications of $\rho_{{\sf N}^{\prime}}$, and allow for coarse grainings of the “purifying” party. Defining $\llbracket{\sf N}\rrbracket=\{0,1,\ldots,{\sf N}\}$, where we added zero to account for the purifier, such a coarse graining can then be specified by a surjective function
@@ IV.1 Subsystem coarse grainings
which specifies, for each party $\ell^{\prime}\in\llbracket{\sf N}^{\prime}\rrbracket$, which of the coarse gained $\llbracket{\sf N}\rrbracket$ parties it belongs to.
Given a density matrix $\rho_{{\sf N}^{\prime}}$ with entropy vector $\vec{{\sf S}}^{\prime}$, and a coarse graining $f$, the components of the entropy vector of the same density matrix after the coarse graining are then given by
${\sf S}^{\prime}_{\mathscr{I}}={\sf S}_{f^{-1}(\mathscr{I})},$ (4)
@@
Notice that in general, this map of entropy vectors from $\Sigma_{{\sf N}^{\prime}}^{*}$ to $\Sigma_{{\sf N}}^{*}$, which we denote by $\Psi_{f}$, can formally be extended to vectors in the full $\overline{\Sigma}^{*}_{{\sf N}^{\prime}}$, and can map a vector on the boundary of $\overline{\Sigma}^{*}_{{\sf N}^{\prime}}$ to one strictly in the interior of $\overline{\Sigma}^{*}_{{\sf N}}$. For example, any coarse graining to ${\sf N}=2$ of the ${\sf N}=3$ entropy vector corresponding to the four-party Greenberger-Horne-Zeilinger (GHZ) state gives an entropy vector strictly inside $\overline{\Sigma}^{*}_{2}$ (since no instance of SA is saturated).
@@ IV.2 Inner bounds for $\overline{\Sigma}^{*}_{{\sf N}}$
Using this construction, we can now derive an inner bound for $\overline{\Sigma}^{*}_{{\sf N}}$ as follows. Suppose that $\mathscr{R}_{{\sf N}^{\prime}}$ is known for some ${\sf N}^{\prime}\geq{\sf N}$. We can then consider all possible coarse graining from ${\sf N}^{\prime}$ to ${\sf N}$, mapping all vectors in $\mathscr{R}_{{\sf N}^{\prime}}$ to ${\sf N}$-party entropy vectors, and taking the conical hull. Formally, for any given ${\sf N}$ and ${\sf N}^{\prime}\geq{\sf N}$, we define
@@
It suffices to show that for any ${\sf N}^{\prime}\geq{\sf N}$ and $\vec{\sf R}\in\mathscr{R}_{{\sf N}}$, there exists some $\vec{\sf R}^{\prime}\in\mathscr{R}_{{\sf N}^{\prime}}$ and a coarse graining $f$ from ${\sf N}^{\prime}$ to ${\sf N}$ such that $\Psi_{f}(\vec{\sf R}^{\prime})=\vec{\sf R}$. For the case where $\vec{\sf R}$ is realized by a density matrix $\rho$, we can realize the desired $\vec{\sf R}^{\prime}$ with $\rho^{\prime}=\rho\otimes\ket{{\bf 0}}\!\bra{{\bf 0}}_{[{\sf N}^{\prime}]\setminus[{\sf N}]}$, and choose $f$ such that the additional factors are coarse gained with the purifier for $\rho^{\prime}$. For the general case, we refer the reader to Section 3.4 of He et al. (2023). ∎
@@ V Hypergraph models
The prescription we just described associates to each ${\sf N}$-party hypergraph model a vector in $\mathbb{R}^{\sf{D}}$. The collection of all these vectors (at fixed ${\sf N}$) is again a convex cone  [^3] A conical combination of two entropy vectors realized by hypergraphs is realized by the hypergraph obtained by the disjoint union of the hypergraphs with the weights appropriately rescaled., and we have the following theorem Walter and Witteveen (2021).
@@
While in principle hypergraph models may not be sufficient to solve the EQMIP completely at arbitrary ${\sf N}$ (see §VII), they at least provide a partial solution that will be sufficient for our purposes.
@@
Figure 1: A hypergraph giving an element of $\mathscr{R}_{6}$. The hyperedge (yellow blobs) has weight 2, and edges (blue lines) have weight 1.
@@
Figure 2: A hypergraph giving an element of $\mathscr{R}_{9}$. The hyperedge (yellow blobs) has weight 2, and edges (blue lines) have weight 1.
@@
Figure 3: A hypergraph giving an element of $\mathscr{R}_{8}$. The two hyperedges (yellow and purple blobs) have weight 2, and edges (blue lines) have weights either 1 (when unspecified) or 2.
@@ VI Constructions for small ${\sf N}$
Having reviewed the necessary tools from hypergraph models, we now use this technology to construct inner bounds for $\overline{\Sigma}^{*}_{{\sf N}}$ for ${\sf N}\leq 4$. The logic will be as follows: we will construct hypergraphs that realize certain extreme rays of the SAC which, by 1, are automatically solutions to the EQMIP. Even without a full solution to the EQMIP, we will then show via particular coarse grainings that we can obtain entropy vectors for a smaller number of parties, which give powerful inner bounds.
@@ VI.1 Bounds for ${\sf N}=2$
This case is trivial, but we include it for completeness. The outer bound $\Sigma_{2}$ to $\overline{\Sigma}^{*}_{2}$ is the SAC, and its extreme rays are realized by Bell pairs. We then have
@@ VI.2 Bounds for ${\sf N}=3$
As we mentioned in §II, it was already shown in Pippenger (2003) that $\overline{\Sigma}^{*}_{3}=\Sigma_{3}$. This equivalence follows from the fact that the extreme rays of $\Sigma_{3}$ contain entropy vectors realized by Bell pairs, the four-party GHZ state, and the four-party absolutely maximally entangled state (also known as the four-party perfect tensor) Helwig and Cui (2013). With the only exception of the four-party GHZ, all these states also generate extreme rays of the three-party SAC and are hence elements of $\mathscr{R}_{3}$. Their conical hull is $\Delta_{3}^{3}$, which coincides with the three-party holographic entropy cone Bao et al. (2015).
To show that the sequence in Eq. (6) converges to $\overline{\Sigma}^{*}_{3}$, we need to consider $\Delta_{3}^{3+i}$ for $i\geq 1$. The sets $\mathscr{R}_{4}$ and $\mathscr{R}_{5}$ were already found in Hernández-Cuenca et al. (2019), and it is straightforward to verify that they imply $\Delta_{3}^{3}=\Delta_{3}^{4}=\Delta_{3}^{5}$. Hence, we need $i\geq 3$. It turns out that $i=3$ is sufficient, since one can verify that the entropy vector associated to the hypergraph in Figure 1 is in $\mathscr{R}_{6}$, and under the coarse graining